# Streamlining and standardizing software citations with The Software Citation Station

### v1

@@ Abstract
Astronomy software (1855) – Software documentation (1869) – Open source software (1866)
***tomwagg@uw.edu, fsb2127@columbia.edu
@@ 1 Introduction
Software development has become an essential part of every sub-field of astronomy, but software developers and large software development efforts are not adequately funded or supported by existing structures (National Academies of Sciences, Engineering, and Medicine, 2021). Citations are an important measure in science and astronomy to build upon and properly attribute and credit earlier work, value authorship, motivate (future) funding, enable peer-review, validate and reproduce findings, support collaboration and community, and encourage reuse and building on work of others (Katz et al., 2020). Many tools have been developed to make astronomy citations easier, such as NASA/ADS [^2] https://ui.adsabs.harvard.edu and ArXiv [^3] https://arxiv.org. However, software citations have long been lagging behind. Software is often not cited properly, it is often left to the software author to point readers to the citation mechanism, and software lacks a standardized citation practice (Howison & Bullard, 2016; Niemeyer et al., 2016; Li et al., 2017; Bouquin et al., 2020; Alsudais, 2021; Bouquin et al., 2023).
[…]
@@
Other advancements focus on the review and reproducibility of software accompanying research. An important example is the creation of The Journal of Open Source Software (JOSS [^4] https://joss.theoj.org; Arfon 2020), an open access journal for research software packages that enables peer-review of astronomy software without processing charges or subscription fees. Recently, AAS publishing started a public collaboration with JOSS where authors submitting to one of the AAS journals can also publish a companion software paper in JOSS (Vishniac & Lintott, 2018a). Additionally, AAS publishing started to facilitate ‘living articles’ – articles that are formally reviewed but that can be updated to accommodate changing conditions, upgrades, or new functionality – to make it easier to publish and update papers describing (long-lived) software (Vishniac & Lintott, 2018b). Similarly, Show your Work! [^5] https://show-your.work (Luger et al., 2021) provides a tool to accompany papers with an automatized workflow that creates a self-contained recipe for readers to reproduce results based on accompanying code and software. Similarly, papers with code [^6] https://paperswithcode.com and Figshare [^7] https://figshare.com are websites aimed to increase visibility of software accompanying (ArXiv) papers.
@@
Another important focus has been on creating unique identifiers to ensure citations are standardised and accurately counted. The Astrophysics Source Code Library (ASCL [^8] https://ascl.net) is a website and tool designed with this goal, specifically dedicated to curating and indexing software used in astronomy-based literature by providing a library of astronomy software with unique identifiers. In a similar vein, GitHub has created an integration with Zenodo to make it easier to reference and cite GitHub repositories in academic literature (GitHub, 2024). Software records in places such as ASCL, Zenodo, and JOSS, which provide the software with an unique identifier, will typically also be included in ADS (with citations counted based on the unique identifier), and efforts are on the way to streamline this between software versions and platforms, e.g., through the AAS Asclepias [^9] https://journals.aas.org/news/asclepias_aas233/ project (Nielsen et al., 2018; van de Sandt et al., 2019; Henneken et al., 2022).
@@
Despite these many advancements, many authors still do not include proper software citations. This is because it can be challenging to find the citation for many tools (or users might not be aware that their software has references), and different software users might end up using different key words or references, making it often difficult to identify and quantify contributions of software and give proper credits to all software contributors (see Katz & Chue Hong, 2018; Katz et al., 2019, 2020; van de Sandt et al., 2019; Bouquin et al., 2020, 2023; Druskat et al., 2021, 2024, and references therein). As citing software is becoming increasingly important to support the careers and field in this area, it is important that astronomy increases the practice of software citation in academic work. To this end we have created ‘The Software Citation Station [^10] https://www.tomwagg.com/software-citation-station/, a publicly available website and tool to make it extremely easy to find and add software citations to academic work.
@@ 2.4 Adding new software
We aim for The Software Citation Station to be open to all software packages used in academic work. As such, the site includes a form for submitting new software packages [^11] https://www.tomwagg.com/software-citation-station/?new-software=true, which guides users through the submission process. Figure 2 shows an example of the online form. Each package requires a data entry of the following form: {minted}[]json ”package name”: ”tags”: [], ”logo”: ””, ”language”: ””, ”category”: ””, ”keywords”: [], ”description”: ””, ”link”: ””, ”attribution_link”: ””, ”zenodo_doi”: ””, ”custom_citation”: ””, ”dependencies”: [] , where tags is a list of the BibTeX tags associated with this package, logo is a path to the logo image file, language is the programming language of the software, category is the category of the package (e.g. “population synthesis” or “visualisation”), keywords are a list of relevant keywords for searching, description is a short 1-2 sentence description of the package, link is a URL to the documentation (or code repository if no documentation is available) and attribution_link a link to a website that outlines the preferred citation format of the package. Optionally each package can also have a zenodo_doi, which is the DOI associated with all versions of the package on Zenodo,custom_citation, which is a custom citation string that overwrites the default generated citation (see Section 2.2 for more details) and dependencies, which is a list of package names that should also be cited if this package is cited (see Section 2.3).
@@ 3 Future work
We created The Software Citation Station to make it incredibly easy for astronomers to add citations of commonly-used software to their academic publications. This is in line with recommendations by Bouquin et al. (2023) who mentioned the need to “Create a website that can act as a central source of information on software citation in practice”. [^12] See also CiteSoftware.org, CiteAs.org, and cite.research-software.org which are online guides and tools on how to cite software, but which do not yet create the BibTeX for a specific software package for you as The Software Citation Station does. On the longer term it would be beneficial to integrate The Software Citation Station with other existing efforts to improve citations of software. This includes connecting or integrating our website as part of ArXiv, ADS, and/or ASCL, as well as connecting to citation software like makecite (Price-Whelan et al., 2019), citepy (Barnes, 2021), and duecredit (Halchenko et al., 2024). Specifically, it would be useful to create a consistent library with citation entries that is community-updated to streamline consistency between different software citation tools and improve completeness of the libraries. Ideally such a library could be generated from data provided by the software developer. An important step to achieve this will be for software-developers in astronomy to use consistent citation practices, which could include adding a citation file format file (e.g. CITATION.cff [^13] https://citation-file-format.github.io/, Druskat et al. 2021) to GitHub or other software repositories as well as a CodeMeta file (CodeMeta.json [^14] https://codemeta.github.io/, Jones et al. 2017) and assigning software references a unique identifier (see Katz & Smith 2014; Cosmo et al. 2020, and references therein). Additionally, such software citations should also support version control (e.g. Chen & Davidson, 2020). This work is out of scope of this paper, but would be a useful topic to discuss and pursue in the future (cf. Katz & Chue Hong, 2018; Katz et al., 2019, 2020; van de Sandt et al., 2019; Cosmo et al., 2020; Allen, 2021; Druskat et al., 2022; Bouquin et al., 2023) and examples of recent efforts in this direction are the HERMES (Druskat et al., 2022), CiteAs (Du et al., 2021), and Asclepias (Henneken et al., 2022) projects.
@@ 4 Additional resources
Several organizations are leading software and coding focused academic schools, many of which also include publicly available recordings and lecture notes. Examples include Code/Astro [^17] https://github.com/semaphoreP/codeastro, the LSSTC data science fellowship program [^18] https://lsstdiscoveryalliance.org/programs/data-science-fellowship/, and the La Serena School for data science [^19] http://lssds.aura-astronomy.org/winter_school/. In addition, the Astronomical Data Analysis Software and Systems (ADASS [^20] https://www.adass.org) organizes annual software-focused conferences and workshops in astronomy.
Other software-focused developments in astronomy include several organizations initiating the awarding of software development prizes to recognize and highlight software contributions in the field, such as the ADASS Prize for an Outstanding Contribution to Astronomical Software [^21] https://www.adass.org/softwareprize.html and the NASA software of the year award [^22] https://www.nasa.gov/ames-exploration-technology/awards/. Additionally, there are efforts to promote more stable career opportunities for software focused astronomers including the Software Research Faculty Award [^23] https://www.simonsfoundation.org/grant/scientific-software-research-faculty-award/ from the Simons Foundation. These efforts are important to address the issues beyond citing software, such as supporting and funding the careers of software developers and software engineers in astronomy (see the discussions in Anzt et al., 2020; Bouquin et al., 2023; Carlin et al., 2023; Sochat, 2024). Moreover the development of tools such as CiteLang (Sochat, 2022) move beyond simple software citations to contributions of the dependencies for software. We also point readers to Smith (2022) for an example of a career path as a software engineer in astronomy and a discussion of the challenges.
TW acknowledges support from NASA ATP grant 80NSSC24K0768. FSB is supported by Simons Foundation award 1141468 hosted at Columbia University. This research has made use of NASA’s Astrophysics Data System.

### v2

@@ 1 Introduction
**footnotetext: tomjwagg@gmail.com, fbroekgaarden@ucsd.edu;
both authors contributed equally to the work.
@@
Software development has become an essential part of every sub-field of astronomy, but is often not adequately supported in astronomy (National Academies of Sciences, Engineering, and Medicine, 2021). Citations are an important tool in science and astronomy to build upon and properly attribute and credit earlier work, value authorship, motivate funding, enable peer-review, validate and reproduce findings, support collaboration and community, and encourage reuse and building on work of others (Katz et al., 2020). While there exists a well-developed ecosystem of tools and services to assist with citations to traditional publications, such as NASA/ADS and ArXiv, this infrastructure does not work nearly so well for software citation. As a result software is often not cited properly. A major issue is that software lacks a standardized citation practice, leaving it to the software author to point readers to the citation mechanism (Howison & Bullard, 2016; Niemeyer et al., 2016; Li et al., 2017; Bouquin et al., 2020; Alsudais, 2021; Bouquin et al., 2023).
Some advancements have been made in recent years to promote and standardize software citations in academic journals. This includes the creation of specific recommendations for software citation such as the reports by the Joint Declaration of Data Citation Principles, and the recommendations of the FORCE11 Software Citation Working Group (Martone et al., 2014; Smith et al., 2016; Katz et al., 2020). Several astronomy journals have developed software policies and include software sections or software acknowledgements in their publication templates (e.g., Vishniac & Lintott, 2016; Timmes & Muench, 2019). Examples include those by AAS publishing (AAS Journals, 2024a), and the examples listed on CHORUS (CHORUS, 2023). This has helped to increase software citations, as evidenced by IRAF citations increasing by a factor of ${\sim}150\%$ after AAS started actively recommending authors to include their citations in submitted papers (AAS Journals, 2019).
@@
Other advancements focus on the review and reproducibility of software accompanying research. An important example is the creation of The Journal of Open Source Software (JOSS [^1] https://joss.theoj.org; Arfon 2020), an open access journal for research software packages that enables peer-review of astronomy software without processing charges or subscription fees. Recently, AAS publishing started a public collaboration with JOSS where authors submitting to one of the AAS journals can also publish a companion software paper in JOSS (Vishniac & Lintott, 2018a). Additionally, AAS publishing started to facilitate ‘living articles’ – articles that are formally reviewed but that can be updated to accommodate changing conditions, upgrades, or new functionality – to make it easier to publish and update papers describing (long-lived) software (Vishniac & Lintott, 2018b). Similarly, Show your Work! (Luger et al., 2021) provides a tool to accompany papers with an automatized workflow that creates a self-contained recipe for readers to reproduce results based on accompanying code and software. Similarly, papers with code and Figshare are websites aimed to increase visibility of software accompanying (ArXiv) papers.
@@
Another important focus has been on creating unique identifiers to ensure citations are standardised and accurately counted. The Astrophysics Source Code Library (ASCL) is a website and tool designed with this goal, specifically dedicated to curating and indexing software used in astronomy-based literature by providing a library of astronomy software with unique identifiers. In a similar vein, GitHub has created an integration with Zenodo to make it easier to reference and cite GitHub repositories in academic literature (GitHub, 2024). Software records in places such as ASCL, Zenodo, and JOSS, which provide the software with an unique identifier, will typically also be included in ADS (with citations counted based on the unique identifier), and efforts are on the way to streamline this between software versions and platforms, e.g., through the AAS Asclepias project (Nielsen et al., 2018; van de Sandt et al., 2019; Henneken et al., 2022).
@@
Despite these many advancements, many authors still do not include proper software citations. This is because it can be challenging to find the citation for many tools (or users might not be aware that their software has references), and different software users might end up using different key words or references, making it often difficult to identify and quantify contributions of software and give proper credits to all software contributors (see Katz & Chue Hong, 2018; Katz et al., 2019, 2020; van de Sandt et al., 2019; Bouquin et al., 2020, 2023; Druskat et al., 2021, 2024, and references therein). As citing software is becoming increasingly important to support the careers and field in this area, it is important that astronomy increases the practice of software citation in academic work. To this end we have created ‘The Software Citation Station [^2] https://www.tomwagg.com/software-citation-station/, a publicly available website and tool to make it extremely easy to find and add software citations to academic work. The Software Citation Station is publicly available through Wagg:2024-scs-paper.
@@ 2.4 Adding new software
We aim for The Software Citation Station to be open to all software packages used in academic work. As such, the site includes a form for submitting new software packages, which guides users through the submission process. Figure 2 shows an example of the online form. Each package requires a data entry of the following form:
"packagename": {
"tags": [],
"logo": "",
"language": "",
"category": "",
"keywords": [],
"description": "",
"link": "",
"attribution_link": "",
"zenodo_doi": "",
"custom_citation": "",
"dependencies": []
where tags is a list of the BibTeX tags associated with this package, logo is a path to the logo image file, language is the programming language of the software, category is the category of the package (e.g. “population synthesis” or “visualisation”), keywords are a list of relevant keywords for searching, description is a short 1-2 sentence description of the package, link is a URL to the documentation (or code repository if no documentation is available) and attribution_link a link to a website that outlines the preferred citation format of the package. Optionally each package can also have a zenodo_doi, which is the DOI associated with all versions of the package on Zenodo,custom_citation, which is a custom citation string that overwrites the default generated citation (see Section 2.2 for more details) and dependencies, which is a list of package names that should also be cited if this package is cited (see Section 2.3).
@@ 3 Future work
We created The Software Citation Station to make it incredibly easy for astronomers to add citations of commonly-used software to their academic publications. This is in line with recommendations by Bouquin et al. (2023) who mentioned the need to “Create a website that can act as a central source of information on software citation in practice”. [^3] See also CiteSoftware.org, CiteAs.org, and cite.research-software.org which are online guides and tools on how to cite software, but which do not yet create the BibTeX for a specific software package for you as The Software Citation Station does. On the longer term it would be beneficial to integrate The Software Citation Station with other existing efforts to improve citations of software. This includes connecting or integrating our website as part of ArXiv, ADS, and/or ASCL, as well as connecting to citation software like makecite (Price-Whelan et al., 2019), citepy (Barnes, 2021), and duecredit (Halchenko et al., 2024). Specifically, it would be useful to create a consistent library with citation entries that is community-updated to streamline consistency between different software citation tools and improve completeness of the libraries. Ideally such a library could be generated from data provided by the software developer. An important step to achieve this will be for software-developers in astronomy to use consistent citation practices, which could include adding a citation file format file (e.g. CITATION.cff, Druskat et al. 2021) to GitHub or other software repositories as well as a CodeMeta file (CodeMeta.json, Jones et al. 2017) and assigning software references a unique identifier (see Katz & Smith 2014; Cosmo et al. 2020, and references therein). Additionally, such software citations should also support version control (e.g. Chen & Davidson, 2020). This work is out of scope of this paper, but would be a useful topic to discuss and pursue in the future (cf. Katz & Chue Hong, 2018; Katz et al., 2019, 2020; van de Sandt et al., 2019; Cosmo et al., 2020; Allen, 2021; Druskat et al., 2022; Bouquin et al., 2023) and examples of recent efforts in this direction are the HERMES (Druskat et al., 2022), CiteAs (Du et al., 2021), and Asclepias (Henneken et al., 2022) projects.
@@ 4 Additional resources
Several organizations are leading software and coding focused academic schools, many of which also include publicly available recordings and lecture notes. Examples include Code/Astro, the LSSTC data science fellowship program, and the La Serena School for data science. In addition, the Astronomical Data Analysis Software and Systems (ADASS) organizes annual software-focused conferences and workshops in astronomy.
Other software-focused developments in astronomy include several organizations initiating the awarding of software development prizes to recognize and highlight software contributions in the field, such as the ADASS Prize for an Outstanding Contribution to Astronomical Software and the NASA software of the year award. Additionally, there are efforts to promote more stable career opportunities for software focused astronomers including the Software Research Faculty Award from the Simons Foundation. These efforts are important to address the issues beyond citing software, such as supporting and funding the careers of software developers and software engineers in astronomy (see the discussions in Anzt et al., 2020; Bouquin et al., 2023; Carlin et al., 2023; Sochat, 2024). Moreover the development of tools such as CiteLang (Sochat, 2022) move beyond simple software citations to contributions of the dependencies for software. We also point readers to Smith (2022) for an example of a career path as a software engineer in astronomy and a discussion of the challenges.
5 Data availability and software
The Software Citation Station is publicly available on Zenodo at Wagg:2024-scs-paper. We welcome contributions to the code which can be made through GitHub or the website.
Acknowledgements.
TW acknowledges support from NASA ATP grant 80NSSC24K0768. This work was supported by Simons Foundation award 1141468 hosted at Columbia University. This research has made use of NASA’s Astrophysics Data System. This research is funded by NASA HPOSS grant 80NSSC25K7555 under award number 316592-00001.